You are looking at a one-year loan of $12,000. The interest rate is quoted as 8 percent plus three points. A point on a loan is simply 1 percent (one percentage point) of the loan amount. Quotes similar to this one are common with home mortgages. The interest rate quotation in this example requires the borrower to pay three points to the lender up front and repay the loan later with 8 percent interest.

What rate would you actually be paying here? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Interest rate             %
Answer #1

Actual funds received = 12000*(1-0.03) = 11640

Interest = 8%*12000 = 960

actual Rate paid = 960/11640 * 100

= 8.25%